# H-STAR: LLM-driven Hybrid SQL-Text Adaptive Reasoning on Tables

### v2

@@ A.4 Benefit of using Transposed Table
Table 14 compares H-STAR’s performance using transposed versus original tables for text-based column verification. On TabFact, H-STAR achieves 86.51% accuracy with transposed tables, versus 85.22% with original tables. On WikiTQ, accuracy is 68.62% with transposed tables, compared to 66.59% with original tables.
@@
Table 14: Comparison between H-STAR using a) transposed table and b) original table for text-based column extraction.
Analysis. Transposing tables leads to better performance than using the original table for column verification. The transposed table approach is more effective for complex datasets like WikiTQ, which are harder to reason upon and contain longer tables.
Appendix B Input Table Format
@@
B.1 Symbolic Reasoning
@@
B.2 Textual Reasoning
We convert the tables into linear, sequential text. Continuing with the format of previous textual reasoning methods Ye et al. (2023); Chen (2023); Wang et al. (2023), we adopt the PIPE encoding i.e. plain text separated by ‘|’. Furthermore, we append the table with the caption and the list of columns similar to the format in Section B.1. Example of the input table:
table caption: 2012–13 Exeter City F.C. season
col : name | league | total
row 1: danny coles | 3 | 3
row 4: john o’flynn | 11 | 12
row 8: jamie cureton | 20 | 20
columns: [’name’, ’league’, ’total’]
Appendix C Comparison with ReAct
While our method appears to be similar to ReAct in terms of combining textual reasoning and tool-based methods, there are significant differences between H-STAR and ReAct:
1. Integration of Symbolic and Semantic Reasoning.
H-STAR combines SQL-based and text-based reasoning, while ReAct merges textual reasoning with actions. ReAct uses tools mainly for actions, not reasoning. H-STAR applies SQL for quantitative tasks and text for lexical queries, whereas ReAct focuses more on textual reasoning.
2. Fixed Multi-Stage Approach.
ReAct prompts LLMs to alternate between reasoning and actions, allowing interaction with external sources. H-STAR, on the other hand, splits tabular reasoning into two stages: 1) Table Extraction, and 2) Adaptive Reasoning, blending SQL and text reasoning. This hybrid approach sets H-STAR apart from ReAct’s focus on textual reasoning and interaction with the environment.
3. Table Extraction Approach.
H-STAR uses a "multi-view" approach for table extraction, identifying relevant columns from both original and transposed tables before selecting rows, reducing irrelevant data, and improving focus. In contrast, ReAct does not employ any extraction method.
4. Adaptive Reasoning Strategy.
H-STAR uses SQL to support semantic reasoning for quantitative tasks while relying solely on semantic reasoning for lexical queries. In contrast, ReAct interleaves reasoning and actions without adapting to task type. H-STAR’s hybrid approach, combining SQL logic and text understanding, excels in handling diverse tasks, outperforming ReAct on table reasoning tasks.
@@ Appendix D Implementation Details
We provide the implementation details including the hyperparameters and prompt details for each of the individual steps. H-STAR operates in two key stages: 1) Table Extraction and 2) Adaptive Reasoning. The Table Extraction phase includes column extraction (SQL-based: colsql and text-based: coltext) and row extraction (SQL-based: rowsql and text-based: rowtext). The Adaptive Reasoning phase consists of SQL-based (fsql) and text-based (ftext) reasoning components.
D.1 Hyperparameters
@@ Appendix E Error Analysis: Case Study
Figures 16 and 17 showcase how semantic and symbolic approaches can individually fail for the column and row extraction steps. Figures 19, 18, 20 and 21 illustrate error cases for each individual step within H-STAR.

### v3

@@ A.3 Advantages of using Transposed Table
Table 14 compares the performance of H-STAR using original versus transposed tables for text-based column verification. In TabFact, H-STAR achieves 86.51% accuracy with transposed tables versus 85. 22% with the original tables. On WikiTQ, accuracy is 68.62% with transposed tables, compared to 66. 59% with the original tables.
@@
Table 14: Comparison between H-STAR using (a) transposed table and (b) original table for text-based column extraction.
Analysis. Transposing tables leads to better performance than using the original table for column verification. The transposed table approach is more effective for complex datasets such as WikiTQ, which are harder to reason on and contain longer tables.
Appendix B Comparison with Other Methods
While H-STAR shares similarities with ReAct and E5 in combining textual reasoning with tool-based methods, there are key differences. Like E5, H-STAR employs table extraction to filter irrelevant data and leverages LLMs’ code generation capabilities for table reasoning. However, there are significant differences.
1. Integration of Symbolic and Semantic Reasoning.
H-STAR combines SQL-based and text-based reasoning, while ReAct merges textual reasoning with actions. While ReAct uses tools primarily for actions and E5 applies code generation only for table extraction, both rely solely on semantic approaches for reasoning. In contrast, H-STAR applies SQL for quantitative tasks and text for lexical queries.
2. Fixed Multi-Stage Approach.
ReAct prompts LLMs to alternate between reasoning and actions, allowing interaction with external sources. H-STAR, on the other hand, splits tabular reasoning into two stages: (1) Table Extraction, and (2) Adaptive Reasoning, blending SQL and text reasoning. This hybrid approach sets H-STAR apart from ReAct’s focus on textual reasoning and interaction with the environment.
3. Table Extraction Approach.
H-STAR distinguishes itself by combining symbolic and semantic methods for table extraction, rather than relying exclusively on either approach. Furthermore, H-STAR uses a ‘multi-view’ approach for table extraction, identifying relevant columns from both original and transposed tables before selecting rows, reducing irrelevant data, and improving focus. In contrast, ReAct does not employ any extraction method.
4. Adaptive Reasoning Strategy.
Unlike E5, H-STAR uses SQL to support semantic reasoning for quantitative tasks while relying solely on semantic reasoning for lexical queries. In contrast, ReAct interleaves reasoning and actions without adapting to task type. H-STAR’s hybrid approach, combining SQL logic and text understanding, excels at handling diverse tasks, outperforming ReAct on table reasoning tasks.
Appendix C Input Table Format
@@
C.1 Textual Reasoning
We convert the tables into linear, sequential text. Continuing with the format of previous textual reasoning methods Ye et al. (2023); Chen (2023); Wang et al. (2023), we adopt the PIPE encoding i.e. plain text separated by ‘|’. Furthermore, we append the table with the caption and the list of columns similar to the format in Section C.2. Example of the input table:
table caption: 2012–13 Exeter City F.C. season
col : name | league | total
row 1: danny coles | 3 | 3
row 4: john o’flynn | 11 | 12
row 8: jamie cureton | 20 | 20
columns: [’name’, ’league’, ’total’]
C.2 Symbolic Reasoning
@@ Appendix D Implementation Details
We provide the implementation details including the hyperparameters and prompt details for each of the individual steps. H-STAR operates in two key stages: (1) Table Extraction and (2) Adaptive Reasoning. The Table Extraction phase includes column extraction (SQL-based: colsql and text-based: coltext) and row extraction (SQL-based: rowsql and text-based: rowtext). The Adaptive Reasoning phase consists of SQL-based (fsql) and text-based (ftext) reasoning components.
D.1 Input Prompts
D.2 Hyperparameters
@@ Appendix E Error Analysis: Case Study
Figures 16 and 17 showcase how semantic and symbolic approaches can individually fail for the column and row extraction steps. Figures 18, 19, 20 and 21 illustrate error cases for each individual step within H-STAR.